# AI’s Social Forcefield: Reshaping Distributed Cognition in Human–AI Teams

### v2

@@ 2.3. Group Identity and Social Cohesion
Social cohesion and team identity emerge from this cognitive coherence, as smooth coordination and shared understanding create a sense of mutual predictability and belonging—turning cognitive alignment into social alignment. Language can signals collective identity through the use of function words that require mutual understanding, such as pronouns (Swol et al., 2021; Sievers et al., 2024). Using pronouns only makes sense when speakers share an understanding of who the pronouns refer to. Groups with frequent pronoun use have a higher level of trust in their teammates and strong social cohesion (Tausczik and Pennebaker, 2010). The level of social cohesion and group identity is crucially dependent on who is considered part of the group. Teams often differentiate between in-group and out-group members, which alters how much alignment occurs and how shared mental models develop (Brennan and Hanna, 2009; Liberman et al., 2018). When an agent—such as an AI system—is treated as a teammate, it may facilitate greater alignment and integration of perspectives (Wang et al., 2021), but if it is seen as an outsider, alignment may be weaker or even resisted (Xie, 2023; Mahmud et al., 2023). Whether AI is integrated into the collaborative fabric of a team or remains at its margins is critical to understand its impact on the overall social fabric of the team (Groom et al., 2011; Nass and Moon, 2000).
In our framework, social cohesion and team identity represent macro-level manifestations of cognitive alignment. We measure them through surface-level linguistic signals—inclusive pronoun use—an approach that is theoretically grounded in the view that language both reflects and enacts shared cognitive and social structures (Hutchins, 1995; Clark, 1996). A substantial body of work demonstrates that function-word choices, particularly pronouns, provide reliable markers of interpersonal alignment, collective identity, and cohesion (Tausczik and Pennebaker, 2010; Gonzales et al., 2010). From a distributed cognition perspective, such linguistic traces reveal the degree to which a team members inhabit a shared representational space—a precondition for successful coordination and collective agency (Pickering and Garrod, 2014; Hutchins, 1995; Searle, 1995).
@@ 3. Related Work
Our research is most closely related to work on (a) human-AI teaming and (b) alignment on shared language between humans and artificial systems.
@@ 3.1. Human-AI Teaming
The growing importance of teamwork (Wuchty et al., 2007) has converged with rapid advances in artificial intelligence to fuel the development of collaborative human–AI teams (Malone and Bernstein, 2015; O’Neill et al., 2020). The vision is that such teams can leverage the complementary strengths of humans and AI, thereby transcending the limitations of human-only groups (Brynjolfsson et al., 2018; Vaccaro et al., 2024).
@@ 3.1.1. From Task Performance to Cognitive Alignment
Most research on human-AI interaction has emphasized performance outcomes, focusing on how AI systems automate processes, augment human decision-making, or generate novel solutions (e.g., Dell’Acqua et al., 2023; Chen and Chan, 2024; Meincke et al., 2024). Experimental setups often compare human-only performance to human-AI dyad performance, with measures such as speed, accuracy, and output quality. For example, studies in medical imaging show that AI can improve diagnostic accuracy when combined with radiologists’ judgments (e.g., Agarwal et al., 2023; Zając et al., 2025), and that AI collaboration can enhance decision-making performance (e.g., Fügener et al., 2022). However, results are inconsistent: some studies report performance gains, while others find disruption or reduced effectiveness (e.g., Autor, 2014; Fügener et al., 2021; Dell’Acqua et al., 2023). This variability highlights the need to look beyond outcomes and investigate the mechanisms that shape collaboration, such as collective attention and cognitive alignment.
Current AI systems are usually optimized for task success rather than for supporting group dynamics like shared focus of attention or collective cognition. This narrow focus can inadvertently disrupt multi-human collaboration, for example, by weakening a team’s ability to sustain collective attention. Research has made significant advances in shedding light on the important of alignment between humans understanding the AI and the AI understanding humans (Bansal et al., 2019a, b; Paleja, 2022; Alipour et al., 2021; Bansal et al., 2019a; Fügener et al., 2022). Much less is known about the indirect “spillover” from AI to human-human dynamics beyond the direct interaction with the AI.
3.1.2. Human-AI Alignment and Ethical Concerns
An emerging body of work explores how humans adapt linguistically and cognitively to AI. Spillner and Wenig (2021) showed that users exhibit lexical alignment with chatbots in digital companion settings, while Li and Song (2025) provided a framework for bidirectional alignment, emphasizing both human-to-AI and AI-to-human processes. However, this work generally focuses on dyadic interactions, leaving open the question of how AI-generated language influences dynamics in larger human groups and beyond the direct interaction with (or presence of) the AI system.
At the same time, researchers have raised ethical concerns about over-adaptation.  van der Rijt et al. (2025) caution that AI mimicking human language too closely can blur accountability and reinforce social biases. Relatedly, work on conversational agents in service encounters demonstrates how AI can subtly shape human communication practices (Van Pinxteren et al., 2020). These findings point to the broader social consequences of AI alignment, highlighting the importance of studying how AI reshapes collaboration beyond immediate task utility, and point to crucial ethical concerns of that arise when humans adapt to AI and/or AI is explicitly designed to steer such adaptation one way or another.
@@ 3.1.3. Toward Multi-Member Dynamics
Other studies have begun to explore how AI systems can enhance group dynamics more directly. For example, shared control between humans and robots allows teammates to combine complementary strengths, improving decision quality and coordination (Music and Hirche, 2017). Similarly, AI agents that scaffold meta-cognitive processes, such as mapping “who knows what” in a group, can strengthen shared mental models and collaboration (Westby and Riedl, 2023). These designs highlight the potential of AI to support not just task outcomes but also the cognitive and social alignment that underpins effective teamwork (Haindl et al., 2022; Bendell et al., 2021; Stowers et al., 2021; Ezer et al., 2019).
Yet despite these advances, most studies remain focused on human-AI dyads (e.g., Bansal et al., 2019a, b; Wang et al., 2019; Fügener et al., 2022; Alipour et al., 2021), leaving fundamental questions about how AI participation alters multi-member team dynamics (Dennis et al., 2023). Collective attention is not simply the sum of individual attention but emerges from multi-agent coordination among teammates and AI (Demir et al., 2019; Harrison et al., 2003). Prior research suggests these dynamics can be unpredictable: if one person adapts smoothly to the AI while another resists, overall team coordination may falter (Balachandar et al., 2019; Reichel et al., 2018).
Guingrich and Graziano (2024) hypothesized that AI-generated language could spill over into human–human collaboration even after the AI is removed, but provided no empirical evidence. Our work directly investigates this proposition, showing how linguistic entrainment to AI reshapes communication within human teams and leaves durable traces in group interaction.
@@ 3.2. Alignment on Shared Language in Human-AI Interaction
When people talk, they co-construct shared terminology to maintain common ground (Branigan et al., 2011; Garrod and Anderson, 1987). With AI, the dynamics differ: users often adapt to the system’s vocabulary to minimize errors, especially if they perceive the AI as less capable of adapting to them (Branigan et al., 2011; Hu et al., 2014; Weatherholtz et al., 2014). Evidence is mixed: some studies suggest stronger alignment with AI than with human partners (Branigan et al., 2011), while others find the opposite (Wudarczyk et al., 2021; Brandstetter et al., 2017). Importantly, alignment with AI sometimes persists beyond the interaction, influencing subsequent human linguistic choices (Brandstetter et al., 2017). Still, most of this work focuses on dyads, leaving unanswered how alignment unfolds in mixed human-AI groups beyond the direct interaction with AI and its broader consequences beyond just language alignment. Gingrich (Guingrich and Graziano, 2024) speculates about AI spillover into human–human teamwork but offers no empirical demonstration. Our study addresses this gap by showing how alignment with AI can leave lasting linguistic and cognitive traces in human group interactions.
3.3. Bridging to Our Research Questions
@@
This review or related work highlights three primary gaps. First, research on human-AI collaboration has focused too narrowly on task outcomes, neglecting the cognitive and social processes that underpin teamwork. Second, existing studies of human-AI alignment are limited to dyadic interactions (one-human-one-AI) and situations of direct AI interaction, leaving open questions about how AI spills over to multi-member teams and situations in which the AI may not be present. Third, foundational work in human communication shows that alignment and entrainment are central to cognitive alignment, suggesting that AI could significantly reshape these processes. Past work has focused on isolated effects without exploring a broader framework of scaffolded effects across multiple layer of alignment. Building on these insights, our work investigates whether and how AI affects cognitive alignment across different layers in human-human interaction, whether alignment with AI language persists when AI is no longer present, and hos it cascades into higher-order cognitive processes including shared attention, mental models, and group identity and cohesion.
4. Study 1 - Team-Based Collaborative Problem-Solving

### v3

@@ 2.3. Group Identity and Social Cohesion
Social cohesion and group identity represent the affective dimension of collaborative alignment (Pickering and Garrod, 2004a; Clark and Wilkes-Gibbs, 1986). When groups align across channels—linguistic, attentional, and representational—they experience cognitive fluency and affective coherence: the feeling that “we understand each other” (Hutchins, 1995; Echterhoff et al., 2009). Convergence itself signals affiliation with the group (Coupland et al., 1991). Shared language, coordinated attention, and aligned mental models thus co-occur with relational bonds and a sense of “we-ness” (Tomasello, 2009; Shteynberg, 2015; Hinds et al., 2014). Crucially, social cohesion depends on who is considered part of the group. Teams differentiate between in-group and out-group members, and this boundary shapes how much alignment occurs (Brennan and Hanna, 2009; Liberman et al., 2018). When an agent—such as an AI system—is treated as a teammate, it may facilitate greater alignment and integration of perspectives (Wang et al., 2021), but if it is seen as an outsider, alignment may be weaker or even resisted (Xie, 2023; Mahmud et al., 2023). Whether AI is considered an insider or outsider is critical to understand its impact on the overall social fabric of the human team members (Groom et al., 2011; Nass and Moon, 2000).
Our operationalization of social cohesion builds on the idea of the dual role of language: language is not only a constructor of group identity but also serves as an index of it (Echterhoff et al., 2009; Hutchins, 1995; Clark, 1996). In particular, language can signal collective identity through the use of function words that require mutual understanding, such as pronouns (Swol et al., 2021; Sievers et al., 2024). Groups with frequent pronoun use have a higher level of trust in their teammates and strong social cohesion (Tausczik and Pennebaker, 2010). We consequently operationalize our measure of alignment on the relational layer through groups’ use of pronouns (Tausczik and Pennebaker, 2010). A substantial body of work has used function-word choices—particularly pronouns—as indicators of interpersonal alignment, collective identity, and cohesion (Tausczik and Pennebaker, 2010; Gonzales et al., 2010; Pickering and Garrod, 2014; Hutchins, 1995; Searle, 1995). We supplement pronoun use as linguistic proxy with survey-based measures of group identity and cohesion to explore the extent to which AI instantiates these patterns.
In summary, we adopt a framework in which AI-induced spillover manifests across parallel channels of socio-cognitive alignment—linguistic, cognitive, and affective—spanning processes that range from automatic to deliberative to social. This breadth is theoretically meaningful: AI spillover could plausibly be confined to automatic linguistic processes alone, without affecting the full spectrum of collaborative cognition.
@@ 3. Related Work
Our research is most closely related to work on (a) human-AI teaming, and (b) alignment on shared language in human-AI teams.
@@ 3.1. Human-AI Teaming
The growing importance of teamwork (Wuchty et al., 2007) has converged with rapid advances in artificial intelligence to fuel the development of collaborative human-AI teams (Malone and Bernstein, 2015; O’Neill et al., 2020). The vision is that such teams can leverage the complementary strengths of humans and AI, thereby transcending the limitations of human-only groups (Brynjolfsson et al., 2018; Vaccaro et al., 2024; Riedl and Weidmann, 2025).
@@ 3.1.1. From Task Performance to Cognitive Alignment
Most research on human-AI interaction has emphasized performance outcomes, focusing on how AI systems automate processes, augment human decision-making, or generate novel solutions (e.g., Dell’Acqua et al., 2026; Chen and Chan, 2024; Meincke et al., 2024). Experimental setups often compare human-only performance to human-AI dyad performance, with measures such as speed, accuracy, and output quality (Vaccaro et al., 2024). For example, studies in medical imaging show that AI can improve diagnostic accuracy when combined with radiologists’ judgments (e.g., Agarwal et al., 2023; Zając et al., 2025), and that AI collaboration can enhance decision-making performance (e.g., Fügener et al., 2022). However, results are often inconsistent and true synergy can be difficult to achieve: some studies report performance gains, while others find disruption or reduced effectiveness (e.g., Autor, 2014; Fügener et al., 2021; Dell’Acqua et al., 2026; Vaccaro et al., 2024). This variability highlights the need to look beyond performance as a singular outcome and investigate the mechanisms that shape collaboration processes overall, such as collective attention, shared mental models, and social cohesion (Riedl and Weidmann, 2025; Kelley et al., 2025). Several studies have started to shed light on different aspects of cognitive alignment (Bansal et al., 2019a, b; Paleja, 2022; Alipour et al., 2021; Bansal et al., 2019a; Fügener et al., 2022); most notably Schelble et al. (Schelble et al., 2022) studied team cognition, performance, and trust. Deeper understanding of how AI exposure affects alignment across different channels could help explain why performance benefits sometimes fail to materialize.
@@ 3.1.2. Toward Multi-Member Dynamics
Most studies of human-AI teaming remain focused on human-AI dyads (e.g., Bansal et al., 2019a, b; Wang et al., 2019; Fügener et al., 2022; Alipour et al., 2021), leaving fundamental questions about how AI participation alters multi-member team dynamics (Dennis et al., 2023). Collective cognition is not simply the sum of individual cognition but emerges from multi-agent coordination among human teammates and AI (Demir et al., 2019; Harrison et al., 2003). Prior research suggests these dynamics can be unpredictable: if one person adapts smoothly to the AI while another resists, overall team coordination may falter (Balachandar et al., 2019; Reichel et al., 2018). Similarly, in-group and out-group dynamics are inherently group phenomena that go beyond the human-AI dyad of direct interaction. Few studies have explored AI’s effect on human-human interaction beyond the direct interaction with the AI itself.
@@ 3.2. Alignment on Shared Language in Human-AI Interaction
An emerging body of work explores how humans adapt linguistically to AI. Spillner and Wenig (2021) showed that users exhibit lexical alignment with chatbots in digital companion settings, while Li and Song (2025) provided a framework for bidirectional alignment, emphasizing both human-to-AI and AI-to-human processes. However, this work generally focuses on dyadic interactions, leaving open the question of how AI-generated language influences dynamics in larger human groups and beyond the direct interaction with (or presence of) the AI system. Guingrich and Graziano (2024) hypothesized that AI-generated language could spill over into human–human collaboration even after the AI is removed, but provided no empirical evidence. Our work directly investigates this proposition, showing how linguistic spillover from AI reshapes communication within human teams and leaves durable traces in group interaction. Brandstetter et al. (Brandstetter et al., 2017) provided insights on linguistic spillover from interaction with robots to subsequent word choices. We expand this work by focusing on word choices during naturalistic language production going beyond naming symbols in a forced-choice task. Furthermore, we introduce precise controls for baseline effects (such as lexical priming tendency and anchoring bias). That is, we capture representational spillover in new contexts, new utterances, new pragmatic roles, not just referential replacement.
@@
When people talk, they co-construct shared terminology to maintain common ground (Branigan et al., 2011; Garrod and Anderson, 1987). With AI, the dynamics are more complicated and empirical evidence is mixed. Some studies suggest stronger alignment with AI than with human partners (Branigan et al., 2011)—especially if they perceive the AI as less capable of adapting to them (Branigan et al., 2011; Hu et al., 2014; Weatherholtz et al., 2014)—while others find the opposite (Wudarczyk et al., 2021; Brandstetter et al., 2017). Still, most of this work focuses on dyads, leaving unanswered how alignment unfolds in mixed human-AI groups beyond the direct interaction with AI and its broader consequences beyond just language alignment. Our study addresses this gap by showing how spillover from AI can leave lasting linguistic and cognitive traces in human-human interactions.
3.3. Summary of Research Gaps
@@
This review or related work highlights three primary gaps. First, existing studies of human-AI alignment are often limited to dyadic interactions (one-human-one-AI) and situations of direct AI interaction, leaving open questions about how AI spills over to multi-member teams and situations in which the AI may not be present. Second, research on human-AI collaboration has focused too narrowly on performance as task outcome, neglecting the cognitive and social processes that underpin teamwork, thus making it difficult to explain situations in which AI fails to deliver performance improvements. Third, foundational work in human communication shows that alignment and entrainment are central to cognitive alignment, suggesting that AI could significantly reshape alignment on different cognitive and social levels. Our work addresses all three gaps: we test for spillover from human-AI to human-human interaction, measure effects across multiple parallel channels of collaborative alignment, and trace these effects beyond the boundaries of active AI engagement.
4. Study 1 - Customer Service Encounters
To establish proof for the existence of causal linguistic spillover, we designed a preregistered [^2] https://osf.io/quj2d/?view_only=9299fa86594b487c9c356e1a32dbe6f7 and IRB approved study. This study uses a minimal design with a state-of-the-art interactive chatbot. It directly tests whether subtle variations in an AI assistant’s framing and language can leave a cognitive and linguistic “fingerprint” that is detectable when people later communicate with other humans—even when the AI is no longer present. The study design randomly manipulates the AI’s system prompt—affecting subtle word choices while keeping the overall task and interaction dynamics constant—and measures how much subjects spontaneously reuse AI words in a face-to-face debriefing interview. This experimental manipulation of linguistic exposure thus isolates the causal pathway through which AI priming reshapes temporally disjoint human language and thought beyond the direct interaction with the AI itself. The only systematic difference between treatments is the AI’s linguistic framing, which allows us to isolate spillover specific to the AI’s style rather than generic entrainment based on common ground or shared tasks.
4.1. Study Design and Task
To achieve high ecological validity, we designed a study around a realistic task: AI assisting customer support agents who perform text-based chat support to resolve customer complaints and technical problems. AI is already widely used for such tasks (BCG, 2024). One prominent field study in a Fortune 500 company showed that AI is highly effective supporting customer service agents, reporting an 14% average increase in productivity (Brynjolfsson et al., 2025). Our study had two phases. First, users interacted with a state-of-the-art chatbot (ChatGPT 4o) (Hurst et al., 2024) to write answers to five customer service complaints across different domains (including technical trouble shooting, online shopping assistance, and reservation booking; one example service complaint was “Every time I try to watch a movie, it buffers endlessly”). The AI was a fully functional ChatGPT 4o assistant: subjects could interact with it like with any chatbot system, write any prompt they liked, for as many rounds as they wanted. Subjects interacted with the AI on a study-provided laptop computer. Users were randomly assigned to one of two AI conditions. Through the system prompt we instructed ChatGPT to either suggest customer service responses that emphasized warmth and empathy or formal adherence to company guidelines (see full system prompts in Appendix A1). The system prompt manipulation subtly changed which words subjects were exposed to by the AI. We term these exposure words and they provide the key experimental manipulation and basis for our analysis. We recorded all user-written prompts and ChatGPT responses (i.e., exposure words), as well as the ultimate response to the customer complaint through a custom user interface implemented via Qualtrics. Each conversation started with ChatGPT offering a suggested draft response to the customer complaint. Subjects wrote on average one additional prompt per customer complaint to refine the initial draft response generated by the chatbot ($SD=1$). The typical interaction was two prompts long. We find no significant different in engagement between the two treatment conditions (two sample t-test, $t=0.843$, $df=92.988$, $p=0.401$).
After the interaction with the AI, subjects participated in an in-person, face-to-face debriefing interview with an experimenter that lasted about five minutes (SD = 108 seconds). The debriefing interview questions were exclusively focused on the tasks—not the AI. To capture not only how the AI shapes how people talk, but also how they imagine acting in the future—their cognitive-behavioral approach to the task—questions were specifically designed to elicit subject’s process, goal, or approach to the task. For example, one question from the interview guide was “If you were working on a real customer service team, what kind of approach would you suggest the team agree on for answering complaints?” A common follow-up question explicitly elicited subject’s approach to the task and behavioral intention “What’s the first thing you usually focus on [when answering customer service complaints]?” In this way, our measure of overlap captures not only linguistic spillover but also how the AI interaction influences how people think about the task itself, thus capturing important aspects of cognitive alignment through which groups sustain a common frame of reference for joint action (Shteynberg, 2015; Clark, 1996; Clark and Brennan, 1991; Tomasello, 2009; Hutchins, 1995).
Interviews were explicitly framed around team-level constructs, asking subjects to describe how a team they might be part of would approach the task (see quote above). This is a well-established technique to study team-level constructs when collecting data from individuals (Mathieu et al., 2022) because “we” narratives are not only expressions of shared language but reflect mutual belief and group identity (Clark, 1996; Gallagher and Tollefsen, 2019). This framing was highly effective: 35% of utterances included the words “we” or “team” (example response from one subject: “[customers] should be let know that we are working on the problem and we resolve it as soon as possible and the team should be assigned as soon as possible to work on it, yeah”—emphasis ours). Having established that subjects engaged with the team-level frame, the AI spillover we detect in these responses can be interpreted as speaking to how AI exposure shapes the way people reason about collective task execution—their sense of how “we as a team” plan to act.
Experimenters were blind to the treatment condition and exposure words that subjects experienced during their task work. They were instructed to maintain a neutral, non-judgmental tone. They were also instructed not to discuss the AI’s language explicitly or reveal the goal of the study. Instead, the debriefing was focus on eliciting detailed verbal descriptions of the task the participant worked on and how they thought about it. We recorded the debriefing interview with wireless lapel microphones via smartphones or laptops and used Amazon Web Services Transcribe to transcribe interviews with automated speaker segmentation. We term the words spoken by subjects during the debrief interview spoken words. Overlap between exposure words and spoken words capture the extent to which the influence from the AI assistant spills over to natural human-human conversations and persists even when AI is no longer an active participant of the interaction.
This is a particularly strong and realistic research design. It mirrors a workplace setting in which individuals interact with AI alone during solo work yet they are embedded in a broader human-human team—a social context in which they develop and discuss shared norms of their task, instruct junior team members how to perform the task, and share behavioral strategies they find effective when completing their task. All of these social interactions can be influenced by individual’s experiences and exposure to AI. This tightly controlled design with a realistic, fully functional chatbot, complements the design of Study 2 in which team language was jointly negotiated, measured shared term use across participants, and included measures shared model alignment. With this experimental design in place, how do we analyze this data?
4.2. Permutation Test
The analysis needs to account for baseline alignment that might arise from common ground and shared tasks alone. For this, we quantify AI induced spillover by comparing the overlap between exposure words subjects saw during their chatbot interaction and the spoken words they spontaneously used later during human-human interaction, against a set of counterfactual exposure words from the other treatment condition. This approach does not test against a “no spillover at all” baseline. Instead, it considers a counterfactual spillover baseline which accounts for expected spillover from the tasks itself (such as the specific customer complaints), common ground, exposure to the same task instructions, and having any conversation with an AI assistant. This permutation approach efficiently accounts for naturally emerging overlap from the task structure and common ground alone. The only systematic difference between the actual exposure and the counterfactual is the AI’s linguistic framing (empathic vs. formal), which allows us to isolate spillover specific to the AI’s style rather than generic entrainment. For example, the word regarding appears in 32 exposure sets of the formal condition, but only in three of the empathic bot. Conversely, the word sorry appears in 35 exposure sets of the empathic bot condition but only three times in the formal bot condition. Yet the words response and please are common in both conditions and thus capture baseline entrainment without differentiating between the two treatment conditions (Table A1 shows example exposure words across treatments). In this way, the permutation test captures spillover of specific AI-styled phrasings that are different between the two treatments, not just shared task vocabulary.
We perform a permutation test that contrast the true overlap with an equally likely counterfactual overlap on the utterance level. Specifically, for each utterance, we constructed a matched counterfactual: the same utterance, but with exposure words drawn from a randomly sampled participant in the opposite treatment condition. The analysis then compares observed vs. counterfactual counts of spillover words for the same utterance. If AI exposure has no effect on subjects spoken words, then treatment assignments are exchangeable, and the observed spillover should look identical compared to the permuted null distribution of spillover. Formally, we test the null hypothesis: AI exposure has no spillover on spoken words beyond task-based baseline spillover. This counterfactual construction provides a rigorous within-utterance (some utterances are longer than others), within-participant (some individuals speak a lot; or are exposed to more words by engaging in more AI turns) control that makes the test more sensitive and interpretable than treating the raw counts in isolation. This analysis isolates the unique spillover signal (observed-counterfactual differences), not just reuse of AI language overall. We perform B=B=2,000 permutation shuffles: each iteration draws one counterfactual set of exposure words for each user. We repeat the analysis using both raw overlap counts, and after removing stopwords and stemming and find similar results (we report the stemmed results).
4.3. Participants
We recruited 97 subjects on campus at a major private US university on the East Coast. Subjects were compensated with a \$10 Amazon gift card. Average age was 24 [18-34], self-reported gender was 46% female, 78% were students, 13% employed, 7% unemployed, 59% held a bachelor’s degree, 30% a master’s, 9% reported some college education.
4.4. Results
We find significant causal spillover of specific AI styled language to later human-human interaction when the AI is no longer present. Not surprisingly, we find strong baseline spillover of shared task vocabulary with an average per-utterance overlap of 2.79 words with counterfactual exposure sets. However, overlap with the true exposure set is 3.00 words, a significantly higher overlap ($p=0.0185$, Figure A1). That is, we find observed spillover that is significantly higher than the expected baseline spillover of shared task vocabulary alone. As robustness test, we repeat the analysis but remove utterances in which the subject explicitly referred to the AI. This analysis captures automatic, naturalistic adoption of AI language—AI words slipping into natural speech—beyond direct rephrasing of AI suggested terms (e.g., we remove utterances like “I feel like the AI was, um, […] I want you to show empathy”). Results are consistent ($p=0.021$, Figure A1). As further robustness test, we replace the raw counts of overlap as test statistic with the coefficient from a Poisson regression which allows us to use additional controls and user-level fixed effects (see Eq. 1). This modeling approach controls for opportunity (utterance length), eliminates all person-level confounds such as baseline tendency to reuse any words (i.e., anchoring bias), engagement level, attention paid to the AI, and controls for utterance-specific factors such as topic, emotion, length, difficulty, and timing. We again find significant increase in AI used words ($\beta=0.151$, 95% CI [0.013 - 0.308], $p=0.025$ based on within–between variance decomposition standard error). That is, exposure to AI words increases the likelihood of using those words in subsequent speech by 16% compared to counterfactual words. For a typical 20-word utterance, this corresponds to roughly one additional word spilling over from the AI to later human-human communication—a subtle but systematic spillover of treatment condition specific AI language. The Appendix reports additional analyses mentioned in the study’s pre-registration which provide additional suggestive evidence of AI’s impact on user language.
5. Study 2 - Team-Based Collaborative Problem-Solving
Having established that AI-specific linguistic spillover exists, Study 2 characterizes its breadth. We move to a dynamic, ecologically complex team setting and measure effects across parallel channels: linguistic, collective attention, shared mental models, and social cohesion. We recruited 20 teams of three or four participants (69 subjects total) to work together on a collaborative problem-solving task via video conferencing. The IRB approved study is designed as a randomized controlled trial with a 2 $\times$ 2 between-subjects design: We manipulated the information provided by the AI (helpful vs. unhelpful) and voice (human-sounding vs. robotic-sounding) of the AI assistant. The experimental design with precisely controlled AI interventions again allows us to causally identify AI induced language spillover (controlling for baseline spillover from shared task vocabulary) and assess its impact on collective attention, shared mental models, and social cohesion.